OpenStudy (anonymous):

an object is dropped from rest at a height of 300m velocity after 2 seconds

OpenStudy (anonymous):

Is there more to your question?

OpenStudy (anonymous):

If you are supposed to think of the object as being a particle with negligible weight and friction then the answer would be given by the equation v = u + at where v is final velocity, u is initial velocity, a is acceleration and t is time. acceleration in this case would be the acceleration of free fall (the speed with which the earth attracts us) and time is 2 seconds. this gives v = 0 + 9.81 * 2 v = 19.62